OpenStudy (anonymous):

what is the sum of the first five terms of the geometric sequence in which a 1=10 and r=1/2 express the answer as an improper fraction using the slash key

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

five is not such a large number that you can't just add them up

OpenStudy (anonymous):

Hey

OpenStudy (misty1212):

\[10+5+2.5+1.25+.625\]

OpenStudy (misty1212):

or if you really want a fraction \[10+5+\frac{5}{2}+\frac{5}{4}+\frac{5}{8}\]

OpenStudy (anonymous):

Is that the answer or

OpenStudy (misty1212):

not too hard since \(10+5=15\) and the common denominator for the last three terms is 8

OpenStudy (misty1212):

no dear that is not the answer that is what you have to add !

OpenStudy (anonymous):

Ok

OpenStudy (misty1212):

use a calculator if you don't want to do the arithmetic http://www.wolframalpha.com/input/?i=10%2B5%2B5%2F2%2B5%2F4%2B5%2F8

OpenStudy (anonymous):

19 3/8

OpenStudy (anonymous):

Is that the answer

OpenStudy (misty1212):

yeah as a mixed number

OpenStudy (misty1212):

but they want an improper fraction

OpenStudy (misty1212):

go with \[\frac{155}{8}\]

OpenStudy (anonymous):

Thanks ur really pretty btw

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]